The restaurant at the Hotel Galaxy offers two choices for breakfast: an all-you-can-eat buffet and an a la carte option, where diners can order from the menu. The buffet option has a budgeted meal price of $52. The a la carte option has a budgeted average price of $41 for a meal. The restaurant manager expects that 40 percent of its diners will order the buffet option. The buffet option has a budgeted variable cost of $32 and the a la carte option averages $25 per meal in budgeted variable cost. The manager estimates that 2,600 people will order a meal in any month.

 

For July, the restaurant served a total of 2,400 meals, including 940 buffet options. Total revenues were $49,820 for buffet meals and $62,780 for the a la carte meals.

 

Required:

a. Compute the activity variance for the restaurant for July.

b. Compute the mix and quantity variances for July.
Answer #1

Step 1) 2,600+ 2,400/2= 2,500    

2400- 940= 1,460

940* (52- 32)+ 1,460* (41-25)= 42,160


step 2) 2,500- 1,460= 1,040          2,500- 940= 1,560

2,400* (1,040/2600)* (52-32)+ 2,400* (1,560/2600)* (41-25)= 42,240


step 3) 1,040* (52-32)+ 1,560* (41-25)= 45,760


mix variable 42,160- 42,240= 80 U answer for part B

quantity variable 42,240- 45,760= 3,520 U for part B


part A) activity variance 80+ 3,520= 3,600 U